OpenStudy (anonymous):

A lock has 24 numbers on it. If a 3-number sequence is to be made without repetitions being allowed, how many different sequences are possible?

OpenStudy (lgbasallote):

im thinking use combination \[nCr = \frac{n!}{r!(n-r)!}\] therefore.. \[24C3 = \frac{24!}{3!(24-3)!}\] does that help?